Required information [The following information applies to the questions displayed below.] In 2018, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2020. Information related to the contract is as follows:

2018 2019 2020

Cost incurred during the year $ 2,204,000 $ 3,192,000 $ 2,424,400

Estimated costs to complete as of year-end 5,396,000 2,204,000 0

Billings during the year 2,140,000 3,256,000 4,604,000

Cash collections during the year 1,870,000 3,200,000 4,930,000

Westgate recognizes revenue over time according to percentage of completion

1. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years. (Do not round intermediate calculations. Loss amounts should be indicated with a minus sign.)
Answer #1

Answer

  • Requirement 1

2018

2019

2020

Revenue

$       2,900,000.00

$       4,200,000.00

$          2,900,000.00

Gross Profits (Loss)

$           696,000.00

$       1,008,000.00

$              475,600.00

--Workings

Working

Year end 2018

Year end 2019

Year end 2020

A

Contract Price

$     10,000,000.00

$    10,000,000.00

$        10,000,000.00

B

Cost Incurred to Date

$       2,204,000.00

$       5,396,000.00

$          7,820,400.00

C

Estimated cost yet to be incurred to complete the contract

$       5,396,000.00

$       2,204,000.00

$                                -   

D = B+C

Total Cost

$       7,600,000.00

$       7,600,000.00

$          7,820,400.00

E = (B/D) x 100

% of Completion

29.00%

71.00%

100.00%

F = A x E

Revenue to date

$       2,900,000.00

$       7,100,000.00

$        10,000,000.00

G

Revenue of Previous year

$                            -  

$       2,900,000.00

$          7,100,000.00

H = F - G

Net Revenue this year

$       2,900,000.00

$       4,200,000.00

$          2,900,000.00

I (=B)

Cost to date

$       2,204,000.00

$       5,396,000.00

$          7,820,400.00

J

Cost to date of previous year

$                            -  

$       2,204,000.00

$          5,396,000.00

K = I - J

Net Cost for the year

$       2,204,000.00

$       3,192,000.00

$          2,424,400.00

L = H - K

Gross Profits

$           696,000.00

$       1,008,000.00

$              475,600.00